OpenStudy (anonymous):

Find the sin, cos, tan, cot, sec, csc: sin(pi/2-feta) =3

ganeshie8 (ganeshie8):

check the question once, sin or cos can never be more than 1

OpenStudy (anonymous):

oops sin(pi/2 - feta) = -3/5

ganeshie8 (ganeshie8):

\(\large \sin (\frac{\pi}{2} - \theta) = \frac{-3}{5}\)

OpenStudy (anonymous):

feta? since when have grades schools been using cheese as variables??

ganeshie8 (ganeshie8):

use this :- \(\large \sin (\frac{\pi}{2} - \theta) = \cos \theta \)

ganeshie8 (ganeshie8):

lol feta is cheese is it

OpenStudy (anonymous):

im not sure how...?

ganeshie8 (ganeshie8):

thats just an identity which you need to remember for now once you have \(\cos \theta = -3/5\) , draw the triangle

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

find the other side using pythagorean formula

OpenStudy (anonymous):

sqrt: 21?

ganeshie8 (ganeshie8):

looks wrong, 5^2 = 3^2 + x^2 25 = 9 + x^2 25-9 = x^2 16 = x^2 ? = x

OpenStudy (anonymous):

sqrt: 16 which is 4

OpenStudy (anonymous):

so now i can find the rest

ganeshie8 (ganeshie8):

yes, u have the reference triangle. so u can find the rest :)

ganeshie8 (ganeshie8):

only trick thing is, how do u figure out the signs ? like, for cos, u got -3/5 wat signs u going to put for others ?

OpenStudy (anonymous):

A solution using Mathematica for the calculations is attached.